# Experimental Analysis of Biasing Voltage Generation in Wave-Controlled RIS

### v1

@@ V Reflective Control Using a Single Biasing Frequency
We study the beam- and null-steering capabilities of the wave-controlled RIS with a single biasing frequency, as visualized in Fig. 16. Although using a single biasing frequency does not allow the complete reflective control permitted by multiple BSWs along the transmission line [16], it will be shown that this strategy can be used for reflective control in the following ways: first for creating nulls at broadside which can, for example, enable designs to minimize the radar cross section, and second for single beam steering, with limitations at large angles. Further possibilities for extending this control will also be discussed.
@@
Figure 16: Wave-controlled RIS consisting of two boards, as described in [13]. The top layer contains $M$ pairs of RIS elements arranged in a row along the $x$-axis, each connected to a varactor diode. In the bottom layer, a single BSW propagates along the BTL, generating the biasing voltages $w(x_{m})$, which are sampled at positions spaced by $d_{x}$ in the $x$-direction.
@@ V-A Nulling the Specular Reflection
For many applications, including those described in [21] and [22], it is necessary to reduce the specular reflection of the incident radiation. Maximizing the reflection towards a different direction does not necessarily fully diminish the specular reflection, due to the non-perfect phase gradient along the RIS. The wave-controlled RIS with multiple BSWs described in [16] was able to suppress the specular component by explicitly including it as a constraint in its optimization criteria. The wave-controlled RIS with a single biasing frequency can also aim at reducing specular reflection, but with fewer control parameters.
Fig. 17(a) shows the RIS reflection at $\theta=0^{\circ}$ for the above design due to normal incident illumination against the biasing frequency $f_{\mathrm{b}}$ and the wave amplitude $W_{\mathrm{b}}$ in (4). The dc bias, $W_{0}$, is not considered to be a tunable parameter here and is set to 4 V as this is the minimum biasing voltage required for the varactor specified in this design (see Table I), and increasing the dc bias will reduce the range of controllable phases. From this colormap, it can be seen that for most combinations of frequency and wave amplitude, the specular reflection at $\theta^{\circ}=0$ is non-negligible. However, there are combinations of these parameters that result in significantly reduced specular reflection. The region of reduced reflection begins at 2 MHz with $W_{\mathrm{b}}=12$ V, and is approximately constant at $W_{\mathrm{b}}=4$ V for frequencies larger than 7 MHz, although the magnitude fluctuates with frequency. The minimal reflections occurring for $W_{\mathrm{b}}=4$ V coincide with the approximately linear region of the voltage-to-phase relationship shown in Fig. 4(4(a)).
@@ V-B Single Beam Control Near Broadside
For large angles, achieving this phase gradient across the RIS leads to phase wrapping. However, for small angles where $|\alpha_{M-1}-\alpha_{0}|<360^{\circ}$ for which phase wrapping is not required, this linear phase response along the RIS can be created by a single BSW. This method works best when $L_{\mathrm{left}}$ is small as it allows for maximum phase variation across the elements. This also follows from the discussion in Sec. IV as the length $L_{\mathrm{right}}$ does not impact the voltage pattern seen by element M1M-1 when the BTL is impedance matched to the AWG.
@@
Due to these limitations, a trade-off exists between the length of the array, $L$, and the maximum usable angle for beam steering, $\theta_{\mathrm{p}}$. Larger arrays will have smaller beamwidths, yet will be more limited in beam steering angles. This is shown in Fig. 18 which illustrates two beam steering examples for both the open- and short-circuit terminations for a wave-controlled RIS with either 27 or 60 elements. At the specified angles, optimal $f_{\mathrm{b}}$ and $W_{\mathrm{b}}$ are obtained through a search of the parameter space (see table II), while $\theta_{p}=0^{\circ}$ is maximized when $W_{\mathrm{b}}=0$ for any frequency. All of the patterns shown in the figure maintain the $|\alpha_{M-1}-\alpha_{0}|<360^{\circ}$ requirement since the radiation patterns for larger $\theta_{\mathrm{p}}$ would have rapidly decreasing directivity in the desired direction.
@@ V-C Single Beam Control for Large Angles
Several cases of radiation patterns using four biasing frequencies that are multiples of $f_{\mathrm{b},0}$ are shown in Fig. 19, each leading to different peak-radiation angles. These radiation patterns illustrate the limitations of the technique discussed in this section. The blue line corresponds to the limit of the near-broadside beam steering when $|\alpha_{M-1}-\alpha_{0}|=360^{\circ}$. The orange and purple lines correspond to biasing frequencies that create symmetric standing wave voltages around the center of the BTL ($w(x_{m})=w(x_{M-1-m})$) in (4)) and thus symmetric beam-splitting radiation patterns around $\theta=0\degree$. The line in yellow shows that, when the standing wave voltage is not symmetric and $|\alpha_{M-1}-\alpha_{0}|>360^{\circ}$, the radiation pattern has large sidelobes that reduce the directivity towards $\theta_{\mathrm{p}}$. This figure also compares the radiation patterns for the current design, shown in Fig. 19(a), to those of a lossless design with unit cells that have a linear voltage-to-phase relationship for the carrier signal $f_{\mathrm{c}}$, shown in Fig. 19(b). The linear relationship is calculated by $\alpha_{m}=360^{\circ}*(V-V_{\mathrm{min}}))/(V_{\mathrm{max}}-V_{\mathrm{min}})$ where $V$ is the voltage applied to the unit cell, and $V_{\mathrm{min}}=4$ and $V_{\mathrm{max}}=15$ are the respective minimum and maximum voltages allowed for the design. These results illustrate that the beam steering method can be improved by using wider bandwidth elements [26].
@@ Appendix A Derivation of the Analytical Model of the RIS Unit Cell
This expression can be reformulated in terms of the so-called electric resonance, $\omega_{\mathrm{e}}$, and magnetic resonance, $\omega_{\mathrm{m}}$. At $\omega_{\mathrm{e}}$, the reflection phase is nearly $180\degree$, while when it is near (but not exactly at) $\omega_{\mathrm{m}}$, the reflection phase is approximately $0\degree$. The magnetic resonance also plays a key role in manipulating the phase of the reflection coefficient and enabling its tunability [27, 28, 29]. These two resonances are defined as
@@
Numerical values for the circuit elements $L_{\mathrm{d}}$, $C_{\mathrm{d}}$, and $R_{\mathrm{d}}$ are determined by performing a single full-wave simulation of the RIS unit cell without including the varactor. In our case, we use CST Studio Suite to model the RIS under an orthogonally incident plane wave traveling along the z-z-axis, thereby accounting for mutual coupling among the RIS elements. The model also incorporates copper losses and the relevant dielectric material properties. From this simulation, the scattering parameters are extracted, and the impedance $Z^{\mathrm{FW}}_{\mathrm{eq}}$ is computed from the Z-parameters to determine the values of $\omega_{\mathrm{e}}$ and $\omega_{\mathrm{m}}$. Once $\omega_{\mathrm{e}}$ and $\omega_{\mathrm{m}}$ are determined, the circuit element values can be obtained as follows:

### v2

@@ V Reflective Control Using a Single Biasing Frequency
We study the beam- and null-steering capabilities of the wave-controlled RIS with a single biasing frequency, as visualized in Fig. 16. Although using a single biasing frequency does not allow for the more flexible reflective control permitted by multiple BSWs along the transmission line [16], it will be shown that this strategy can be used for reflective control in the following ways: first for creating nulls at broadside which can, for example, enable designs to minimize the radar cross section, and second for single beam steering, with limitations at large angles. Further possibilities for extending this control will also be discussed.
@@
Figure 16: Wave-controlled RIS consisting of two boards, as described in [13]. The top layer contains $M$ pairs of RIS elements arranged in a row along the $x$ direction, each connected to a varactor diode. In the bottom layer, a single BSW propagates along the BTL, generating the biasing voltages $w(m)$, which are sampled at positions spaced by $d_{x}$ in the $x$ direction.
@@ V-A Nulling the Specular Reflection
For many applications, including those described in [24] and [25], it is necessary to reduce the specular reflection of the incident radiation. Maximizing the reflection towards a different direction does not necessarily fully diminish the specular reflection, due to the non-perfect phase gradient along the RIS. The wave-controlled RIS with multiple BSWs described in [16] was able to suppress the specular component by explicitly including it as a constraint in its optimization criteria. The wave-controlled RIS with a single biasing frequency can also reduce the specular reflection, but with more limited degrees of freedom.
Fig. 17(a) shows the RIS reflection at $\theta=0^{\circ}$ for the above design due to normal incident illumination against the biasing frequency $f_{\mathrm{b}}$ and the wave amplitude $W_{\mathrm{b}}$ in (4). The dc bias, $W_{0}$, is not considered to be a tunable parameter here and is set to 4 V as this is the minimum biasing voltage required for the varactor specified in this design (see Table I), and increasing the dc bias will reduce the range of controllable phases. From this colormap, it can be seen that for most combinations of frequency and wave amplitude, the specular reflection at $\theta=0^{\circ}$ is non-negligible. However, there are combinations of these parameters that result in significantly reduced specular reflection. The region of reduced reflection begins at 2 MHz with $W_{\mathrm{b}}=12$ V, and is approximately constant at $W_{\mathrm{b}}=4$ V for frequencies larger than 7 MHz, although the magnitude fluctuates with frequency. The minimal reflections occurring for $W_{\mathrm{b}}=4$ V coincide with the approximately linear region of the voltage-to-phase relationship shown in Fig. 4(4(a)).
@@ V-B Single Beam Control Near Broadside
For large angles, achieving this phase gradient across the RIS leads to phase wrapping. However, for small angles where $|\alpha_{M-1}-\alpha_{0}|<360^{\circ}$ which do not lead to phase wrapping, this linear phase response along the RIS can be created by a single BSW. This method works best when $L_{\mathrm{left}}$ is small as it allows for maximum phase variation across the elements. This also follows from the discussion in Sec. IV as the length $L_{\mathrm{right}}$ does not impact the voltage pattern seen by element M1M-1 when the BTL is impedance matched to the AWG.
@@
Due to these limitations, a trade-off exists between the length of the array, $L$, and the maximum usable angle for beam steering, $\theta_{\mathrm{p}}$. Larger arrays will have smaller beamwidths but will be more limited in beam steering angles. This is shown in Fig. 18 which illustrates two beam steering examples for both the open- and short-circuit terminations for a wave-controlled RIS with either 27 or 60 elements. At the specified angles, optimal $f_{\mathrm{b}}$ and $W_{\mathrm{b}}$ are obtained through a search of the parameter space (see Table II), while $\theta_{p}=0^{\circ}$ is maximized when $W_{\mathrm{b}}=0$ for any frequency. All of the patterns shown in the figure maintain the $|\alpha_{M-1}-\alpha_{0}|<360^{\circ}$ requirement since the radiation patterns for larger $\theta_{\mathrm{p}}$ would have rapidly decreasing directivity in the desired direction.
@@ V-C Single Beam Control for Large Angles
Several cases of radiation patterns using four biasing frequencies that are multiples of $f_{\mathrm{b},0}$ are shown in Fig. 19, each leading to different peak-radiation angles. These radiation patterns illustrate the limitations of the technique discussed in this section. The blue line corresponds to the limit of the near-broadside beam steering when $|\alpha_{M-1}-\alpha_{0}|=360^{\circ}$. The orange and purple lines correspond to biasing frequencies that create symmetric standing wave voltages around the center of the BTL (w(m)=w(M1m)w(m)=w({M-1-m})) in (4)) and thus symmetric beam-splitting radiation patterns around $\theta=0^{\circ}$. The line in yellow shows that, when the standing wave voltage is not symmetric and $|\alpha_{M-1}-\alpha_{0}|>360^{\circ}$, the radiation pattern has large sidelobes that reduce the directivity towards $\theta_{\mathrm{p}}$. This figure also compares the radiation patterns for the current design, shown in Fig. 19(a), to those of a lossless design with unit cells that have a linear voltage-to-phase relationship for the carrier signal $f_{\mathrm{c}}$, shown in Fig. 19(b). The linear relationship is calculated by $\alpha_{m}=360^{\circ}*(V-V_{\mathrm{min}}))/(V_{\mathrm{max}}-V_{\mathrm{min}})$ where $V$ is the voltage applied to the unit cell, and $V_{\mathrm{min}}=4$ and $V_{\mathrm{max}}=15$ are the respective minimum and maximum voltages allowed for the design. These results illustrate that the beam steering method can be improved by using wider bandwidth elements [29].
@@ Appendix A Derivation of the Analytical Model of the RIS Unit Cell
This expression can be reformulated in terms of the so-called electric resonance $\omega_{\mathrm{e}}$ and magnetic resonance $\omega_{\mathrm{m}}$. At $\omega_{\mathrm{e}}$, the reflection phase is nearly $180^{\circ}$, while when it is near (but not exactly at) $\omega_{\mathrm{m}}$, the reflection phase is approximately $0^{\circ}$. The magnetic resonance also plays a key role in manipulating the phase of the reflection coefficient and enabling its tunability [30, 31, 32]. These two resonances are defined as
@@
The numerical values for the circuit elements $L_{\mathrm{d}}$, $C_{\mathrm{d}}$, and $R_{\mathrm{d}}$ are determined by performing a single full-wave simulation of the RIS unit cell shown in Fig. 2 without including the varactor, the biasing via at the center of the patch, or the coplanar waveguide on the RIS board ground plane (it includes the via to GND and the small rectangular patch connected to it that will connect to the varactor). In our case, we use CST Studio Suite to model the RIS under an orthogonally incident plane wave polarized along $y$ and traveling along the z-z direction, thereby accounting for mutual coupling among the RIS elements (we use periodic boundary conditions in $x$ and $y$). The model also incorporates copper losses and the relevant dielectric material properties. From this simulation, the scattering parameters are extracted, and the impedance $Z^{\mathrm{FW}}_{\mathrm{eq}}$ is computed from the Z-parameters to determine the values of $\omega_{\mathrm{e}}$ and $\omega_{\mathrm{m}}$. For simplicity, $\omega_{\mathrm{e}}$ and $\omega_{\mathrm{m}}$ are approximated by determining when the phase is exactly $180^{\circ}$ and $0^{\circ}$, respectively. Then, the circuit element values are obtained as follows: